OpenStudy (anonymous):

Solve for A: B=5/7(A-11)

OpenStudy (austinl):

\(B=\dfrac{5}{7(A-11)}\) Is this correct?

OpenStudy (anonymous):

\[B=\frac{ 5 }{ 7 }(A-11)\]

OpenStudy (austinl):

Even easier :P

OpenStudy (austinl):

First I would divide each side by \(\dfrac{5}{7}\) \(\dfrac{7B}{5}=A-11\) Then you add 11 to both sides, \(\dfrac{7B}{5}+11=A\) Make sense? :)

OpenStudy (anonymous):

why is the 7 become the numerator?

OpenStudy (anonymous):

\[A=\frac{ 5 }{ 7 }B+11\] so is this the answer?

OpenStudy (austinl):

Because when you divide by a fraction, you are multiplying by the reciprocal. Lets say you want to do this, \(\dfrac{1}{2}\div\dfrac{3}{4}\) To be able to do this operation you need to multiply the half by the reciprocal of three-fourths. \(\dfrac{1}{2}\times\dfrac{4}{3}\) Which then leads us to the answer, \(\dfrac{4}{6}=\dfrac{2}{3}\) Understand?

OpenStudy (anonymous):